OpenStudy (anonymous):

Gas Calculations Help!! Mass of magnesium Strip (grams): 0.032 g Volume of gas collected (mL): 30 mL Barometric pressure (atm): 1.1 atm Room temperature: 22C Vapor pressure of the water (torr): 19.8 torr

OpenStudy (anonymous):

1. Write the balanced equation for the reaction conducted in this lab, including appropriate phase symbols. Mg (s) + 2 H20 (l) --> H2 (g) + Mg (OH)2 (aq) 2. Determine the partial pressure of the hydrogen gas collected in the gas collection tube. 1.1atm = P1 + 0.0395atm 1.1atm – 0.0395atm = 1.0605 3. Calculate the moles of hydrogen gas collected. n = 1.0605atm * 30L / .0821 * 297 = 31.815 / 24.3837 = 1.3 mol 4. If magnesium was the limiting reactant in this lab, calculate the theoretical yield of the gaseous product. Show all steps of your calculation. 5. Determine the percent yield of this reaction, showing all steps of your calculation.

OpenStudy (anonymous):

I don't remember how to find the theoretical yield or the percent yield. And I'm not sure if the other steps are right

OpenStudy (wolfe8):

Theoretical is just the yield from calculation. Percent yield is the actual yield over theoretical yield times 100%

OpenStudy (abb0t):

Why do you ask so many questions at once?